OpenStudy (anonymous):

How many solutions are there to the equation below? 16(x + 2) - 8 = 16x + 24 A. Infinitely many B. 1 C. 0

OpenStudy (anonymous):

that equals, 16 x + 32 - 8 = 16 x +24

OpenStudy (anonymous):

then, you have 16x + 24 = 16x + 24. bring over the 24. you get. 16 x = 16x

OpenStudy (anonymous):

divide by 16, you get x = x.

OpenStudy (anonymous):

so no matter what x you pick, or number for x, the equation will be the same. there are infinitely many solutions. and just to complete it, you get x - x = 0 bcz you bring the x from the right over to the left side.